# A Comprehensive Framework for Semantic Similarity Detection Using Transformer Architectures and Enhanced Ensemble Techniques

### v1

@@ Abstract
Detecting AI-generated text, especially in short-context documents, is difficult because there is not enough context for accurate classification. This paper presents a new teacher-student model that uses domain adaptation and data augmentation to solve these problems. The teacher model, which combines DeBERTa-v3-large and Mamba-790m, learns semantic knowledge through domain-specific fine-tuning. The student model handles short-context text more efficiently. The system uses a Mean Squared Error (MSE) loss function to guide the student’s learning, improving both accuracy and efficiency. Also, data augmentation methods like spelling correction and error injection make the model more robust. Experimental results show that this approach works better than baseline methods, proving its usefulness for real-time AI-generated text detection and other text classification tasks.
@@
AI-generated text detection, teacher-student model, data augmentation, short-context document classification, DeBERTa-v3-large
@@ I Introduction
The rise of AI-generated content, driven by language models like GPT-3 and BERT, has created problems for content moderation and text classification. Detecting machine-generated text is important for many uses, including combating misinformation and verifying academic work. A main problem is detecting AI-generated text in short-context documents, where traditional models often fail due to limited context. Short-context texts, such as tweets, headlines, or excerpts, lack the longer structures that models usually rely on for classification. As a result, even small changes in the text can have a big impact on performance.
To solve this problem, we propose a teacher-student model that combines the strengths of large pre-trained models and short-context text classification. The teacher model, made up of DeBERTa-v3-large and Mamba-790m, is fine-tuned with domain adaptation. This helps it capture more detailed semantic features and better detect AI-generated content. The teacher model helps guide the student model, which is optimized for short-context handling and efficiency. The student model is smaller and more efficient, trained to predict the teacher’s output using a Mean Squared Error (MSE) loss function. This allows it to benefit from the teacher’s knowledge while being fast enough for real-time use.
We also improve the model’s robustness with data augmentation. By applying text changes like spelling correction, character removal, and adding errors, the model learns to handle real-world variations. These augmentations help the model generalize better, making it more accurate in detecting AI-generated text, even in noisy or imperfect conditions. The teacher-student framework, domain adaptation, and data augmentation together create a model that performs well in short-context text classification, where other models struggle.
@@ II Related Work
Detecting AI-generated text has become an important research area because of the growth of large language models (LLMs) and generative adversarial networks (GANs). Yan et al.[1] discuss generative LLMs, focusing on the challenges they present for distinguishing AI-generated text from human-written text. These challenges are also present with GANs, which Gui et al.[2] discuss in terms of their applications for content generation. GANs are particularly used for creating fake text, a problem addressed by Zellers et al.[3] in their study on defending against fake news generated by neural networks.
To improve detection of AI-generated text, Chakraborty et al.[4] review different methods for identifying machine-generated content, looking at how various AI models perform. A major challenge is the domain specificity of text, as models trained on general datasets may not work well in specialized areas. For example, SciBERT, a model for scientific text, has been successful in detecting AI-generated academic papers by recognizing domain-specific language patterns[5]. Dehaerne et al.[6] also explore machine learning for detecting machine-generated code, highlighting difficulties in identifying such content.
Some studies focus on user interactions with AI-generated content. Lu[7] suggests using decision trees and TF-IDF to improve chatbot user satisfaction, which can also help detect AI-generated dialogue. Li[8] looks at how multimodal data can improve product recommendations, a method that could also be used to detect AI-generated content by combining different data sources.
@@
The problem of detecting AI-generated text involves many areas, such as fake news, code, and academic papers. By using domain-specific models, machine learning techniques, and better detection algorithms, progress can be made in identifying AI-generated content.
@@ III Methodology
Semantic similarity between phrases is a critical task in the patent search and examination process. This paper presents a comprehensive study on advanced deep learning models and ensemble strategies for semantic similarity detection. By integrating transformer-based architectures, bidirectional LSTM layers, and novel tricks like adversarial weight perturbation and linear attention pooling, we achieve state-of-the-art performance. Additionally, we introduce dynamic target grouping and fine-tuned ensemble methods to boost diversity and robustness, ensuring superior generalization. The pipline of model is shown in Fig 1.
@@ V Experiment Results
Table I provides a detailed comparison of the proposed models and ablation studies across the evaluation metrics. The changes in model training indicators are shown in Figure 5.
Figure 5: Model indicator change chart.
TABLE I: Performance Comparison and Ablation Studies
@@ VI Conclusion
This study demonstrates the effectiveness of leveraging transformer-based architectures with Bi-LSTM enhancements, adversarial weight perturbation, and dynamic preprocessing strategies for semantic similarity detection in patent documents. The integration of diverse models, combined with linear attention pooling and target shuffling, significantly improves robustness and accuracy. The ensemble strategy achieves state-of-the-art performance across multiple evaluation metrics, setting a robust foundation for practical applications in patent search and examination processes. Future work will explore domain-specific pretraining and advanced augmentation techniques to further enhance model generalization.

### v2

# A Comprehensive Framework for Semantic Similarity Analysis of Human and AI-Generated Text Using Transformer Architectures and Ensemble Techniques

@@ Abstract
The rapid advancement of large language models (LLMs) has made detecting AI-generated text an increasingly critical challenge. Traditional methods often fail to capture the nuanced semantic differences between human and machine-generated content. We therefore propose a novel approach based on semantic similarity analysis, leveraging a multi-layered architecture that combines a pre-trained DeBERTa-v3-large model, Bi-directional LSTMs, and linear attention pooling to capture both local and global semantic patterns. To enhance performance, we employ advanced input and output augmentation techniques such as sector-level context integration and wide output configurations. These techniques enable the model to learn more discriminative features and generalize across diverse domains. Experimental results show that this approach works better than traditional methods, proving its usefulness for AI-generated text detection and other text comparison tasks.
@@
AI-generated text detection, semantic analysis, multi-layer neural network, data augmentation, pre-trained language model
@@ I Introduction
The rise of AI-generated content, driven by language models like ChatGPT, has created problems for content moderation and text classification. Detecting machine-generated text is important for many use cases, including combating misinformation and verifying academic work. Traditional detection methods, which often rely on surface-level features such as syntax and word frequency, struggle to capture the nuanced differences in how humans and machines construct meaning. This study instead proposes a novel approach based on semantic similarity analysis, which focuses on the underlying patterns of semantic relationships rather than surface-level features. Our hypothesis is that while human and AI-generated text can convey similar meanings, they differ fundamentally in how semantic relationships are structured and maintained. These differences, though subtle, can be detected through careful analysis of semantic patterns.
Our approach leverages a pre-trained DeBERTa-v3-large model as the foundation, which provides robust semantic understanding through its disentangled attention mechanism and enhanced mask decoder. This is particularly effective for capturing subtle semantic differences, as DeBERTa’s ability to separately model content and position information allows for more precise analysis of semantic relationships. The model is further enhanced with two layers of Bi-directional LSTM to capture sequential dependencies and long-range semantic patterns, which are crucial for identifying the characteristic evolution of ideas in human writing. A linear attention pooling mechanism is then employed to focus on the most relevant semantic features, reducing noise and improving the model’s ability to distinguish between human and AI-generated patterns. The final output is produced through a fully connected layer, ensuring robust classification.
To further improve performance, we employed several advanced input augmentation techniques such as Electra models pre-trained with Replaced Token Detection (RTD) objectives, sector-level context concatenation, adversarial weight perturbation, and dynamic target shuffling to improve the model’s robustness. We also applied a wide output configuration to allow our model to capture both local and global semantic patterns effectively, making it particularly adept at identifying the subtle differences between human and AI-generated text.
These enhancements help the model generalize better, making it more accurate in distinguishing AI-generated text from those of human origin, and it is proven capable of achieving state-of-the-art performances in our metrics of choice.
@@ II Related Work
Detecting AI-generated text has become an important research area due to the growth of large language models (LLMs) and generative adversarial networks (GANs). Yan et al.[1] discuss generative LLMs, focusing on the challenges they present to distinguish AI-generated text from human-written text. These challenges are also present with GANs, which Gui et al.[2] discuss in terms of their applications for content generation. GANs are particularly used for creating fake text, a problem addressed by Zellers et al.[3] in their study on defending against fake news generated by neural networks.
To improve the detection of AI-generated text, Chakraborty et al.[4] review different methods for identifying machine-generated content, looking at how various AI models perform. A major challenge is the domain specificity of text, as models trained on general datasets may not work well in specialized areas. For example, SciBERT, a model for scientific text, has been successful in detecting AI-generated academic papers by recognizing domain-specific language patterns[5]. Dehaerne et al.[6] also explore machine learning to detect machine-generated code, highlighting the difficulties in identifying such content.
Some studies focus on user interactions with AI-generated content. Lu[7] suggests the use of decision trees and TF-IDF to improve the satisfaction of chatbot users, which can also help detect AI-generated dialogue. Li[8] examines how multimodal data can improve product recommendations, a method that could also be used to detect AI-generated content by combining different data sources.
@@ III Methodology
This section presents a comprehensive framework using deep learning methods and ensemble techniques for semantic similarity detection. By integrating transformer-based architectures, bidirectional LSTM layers, and novel tricks such as Adversarial Weight Perturbation (AWP) and linear attention pooling, we achieve state-of-the-art performance. Additionally, we introduce dynamic target grouping and fine-tuned ensemble methods to boost diversity and robustness, ensuring superior generalization. The pipline of model is shown in Fig 1.
@@ V Experiment Results
Table I provides a detailed view of performance gains across the evaluation metrics in the ablation study. The losses and performance indicator metrics in each training epochs of the final model are shown in Figure 5.
Figure 5: Performance indicator change chart
TABLE I: Ablation Study results
@@ VI Conclusion
This study demonstrates the effectiveness of leveraging transformer-based architectures with Bi-LSTM enhancements, adversarial weight perturbation, and dynamic preprocessing strategies for comparing the semantic similarity between human and AI-generated text. The integration of diverse models, combined with linear attention pooling and target shuffling, significantly improves robustness and accuracy. The ensemble strategy achieves state-of-the-art performance across multiple evaluation metrics, setting a robust foundation for practical applications in patent search and examination processes. Future work will explore domain-specific pretraining and other augmentation techniques to further enhance model generalization.